The perimeter of a rectangle is 74m. The length is 1 m more than twice the width. Find the demensions.
What is the length?
What is the width?
Can someone help?
I would also like to thank EVERYONE who has helped me since I started having to come here for your help you have been Great and I really apprciate learning from you all.

Answer by checkley77(12844)

2L+2W=74
L=2W+1
2(2W+1)+2W=74
4W+2+2W=74
6W=74-2
6W=72
W=72/6
W=12 ANSWER FOR THE WIDTH.
L=2*12=1
L=24+1
W=25 ANSWER FOR THE L3HGTH.
PROOF:
2*12=2*25=74
24+50=74
74=74